A gas has a volume of 65 mL at a temperature of 70 °C. At what temperature (in Kelvin) will the gas has a volume of 43 mL?


1
Expert's answer
2022-01-17T01:29:03-0500

Using Charles law,we have

V1/T1=V2/T2

T2=(V2×T1)/V1

T2=(43×343)/65

T2=226.91K